# Latent Diffusion Prior Enhanced Deep Unfolding for Snapshot Spectral Compressive Imaging

### v1

@@ 4 Experiments
We conduct experiments on both simulation and real HSI datasets. Following the approaches in Meng et al. [37, 36], Huang et al. [23], Cai et al. [5], we select a set of 28 wavelengths ranging from 450-650nm by employing spectral interpolation techniques applied to the HSI data.
@@ 4.1 Experimental Settings
Simulation and Real Datasets: We adopt two widely used HSI datasets, i.e., CAVE [41] and KAIST [11] for simulation experiments. The CAVE dataset comprises 32 HSIs with a spatial size of 512 $\times$ 512. The KAIST dataset includes 30 HSIs with a spatial size of 2704 $\times$ 3376. Following previous works [37, 36, 23, 5], we employ the CAVE dataset as the training set, while 10 scenes from the KAIST dataset are utilized for testing. During the training process, a real mask of size 256 $\times$ 256 pixels is applied. In our real experiment, we utilized the HSI dataset captured by the SD-CASSI system in Meng et al. [37]. The system captures real-world scenes of size 660 $\times$ 714$\times$ 28 with wavelengths spanning from 450 to 650 nm and dispersion of 54 pixels. Implementation Details: For the diffusion settings, the iteration number $T$ of the diffusion is set to 16, and the latent space dimension $N$ is set to 256. For all phases of training, we use the Adam [24] optimizer and set the learning rate to $4\times 10^{-4}$. PSNR and SSIM [57] are utilized as our metrics. Our method is implemented with the PyTorch and trained using NVIDIA RTX3090 GPUs. More details can be seen in the supplementary material (SM).
@@ 4.2 Compare with State-of-the-art
We numerically compare our proposed method with previous methods including the end-to-end networks: DADF-Net [59], CST [4], BIRNAT [10]; the deep unfolding methods: RDLUF-MixS2 [12], PADUT [28], DAUHST [6], DNU [55]; and traditional model-based methods: TwIST [2] and DeSCI [31]. The visual comparisons on both synthetic and real datasets are also conducted with recent state-of-the-art (SOTA) methods.
Synthetic data: The numeric comparisons on synthetic data can be seen in Table 1. Our proposed method surpasses the recent SOTA method RDLUF-MixS2 (+0.44 dB) according to average PSNR and SSIM. Fig. 5 shows the visual reconstruction results. Three wavelengths including striking colors in RGB reference red, yellow, and green are selected to compare. The golden box part in the reference was chosen to calculate and compare the wavelength accuracy. The accuracy metric is the correlation coefficient with the ground truth of the chosen region, i.e. the ‘Corr’ in the curves. According to the ‘Corr’, our method (0.9935) has a more accurate wavelength curve than others. The zoomed part in the figure also demonstrates that our method has edges of bird wings and bird eyes are clearer than others.
Real data: Two scenes of real SD-CASSI measurement reconstruction results are shown in Fig. 6, and two obvious color regions in each scene RGB references are selected to compare. We can see that our method can reconstruct more details such as the light and dark intricacies of the flower.
@@ 5 Ablation Study
In this ablation study, we train our model on the synthetic training data with 3 unfolding stage models. The results are summarized in Table 2, where ‘w/o prior’ denotes that we only use measurement as the latent encoder input instead of clean HSIs, ‘w/o prior, and TT’ denotes that the latent encoder, diffusion, and prior flow in the Trident Transformer are removed. The training process only conducts a simple one-phase end-to-end strategy instead of a two-phase training. The ablation illustrates that with LDM prior assistance, we can achieve better reconstruction results, and our design of the Trident Transformer successfully aggregates three types of information and effectively compensates for some reconstruction defects. Fig. 2 also visualizes the feature map changes before and after prior enhancement. The enhanced features demonstrate increased concentration on significant parts and edges. Moreover, without accurate guidance, the LDM will even harm the reconstruction. We also compare the influence of diffusion time steps in Fig. 1, 16 steps are enough for good reconstruction results. The inference time in Table 2 illustrates that time is still in a reasonable range even using diffusion 16 steps.
@@ 6 Conclusion
In this paper, we introduce a novel spectral reconstruction network that leverages prior knowledge from the latent diffusion model. The network uses a Trident Transformer to effectively combine physics-driven deep unfolding and powerful latent diffusion model. It achieves state-of-the-art performance on both simulated data and real data and has considerable improvements in computational efficiency. This new strategy of using deep unfolding can shed light on other low-level vision and computational imaging tasks.

### v2

@@ 5 Experiments
We conduct experiments on both simulation and real HSI datasets. Following the approaches in [45, 43, 26, 5], we select a set of 28 wavelengths ranging from 450-650nm by employing spectral interpolation applied to the HSI data.
@@ 5.1 Experimental Settings
Simulation and Real Datasets: We adopt two widely used HSI datasets, i.e., CAVE [48] and KAIST [14] for simulation experiments. The CAVE dataset comprises 32 HSIs with a spatial size of 512 $\times$ 512. The KAIST dataset includes 30 HSIs with a spatial size of 2704 $\times$ 3376. Following previous works [45, 43, 26, 5], we only employ the CAVE dataset as the training set both in Phase I and II, while 10 scenes from the KAIST dataset are utilized for testing. During the training process, a real mask of size 256 $\times$ 256 pixels is applied. In our real experiment, we utilized the HSI dataset captured by the SD-CASSI system in [45]. The system captures real-world scenes of size 660 $\times$ 714$\times$ 28 with wavelengths spanning from 450 to 650 nm and dispersion of 54 pixels.
Implementation Details: For the diffusion settings, the iteration number $T$ of the diffusion is set to 16, and the latent space dimension $N$ is set to 16. Training Phase I (train DUN and LE) needs 200 epochs and Phase II (train DU, $\bm{\epsilon}_{\theta}$, and LE’) needs 100 epochs. For all phases of training, we use the Adam [27] optimizer and cosine scheduler. PSNR and SSIM [67] are utilized as our metrics. Our method is implemented with the PyTorch and trained using NVIDIA RTX3090 GPUs. More details can be seen in the supplementary material (SM).
@@ 5.2 Compare with State-of-the-art
We compare our method with previous methods including the end-to-end networks: DADF-Net [69], MST [5], BIRNAT [13]; the deep unfolding methods: RDLUF-MixS2 [15], PADUT [32], DAUHST [6], DNU [65]; the self-supervised method: LRSDN [11]; and traditional model-based method: TwIST [2]. The comparisons are conducted on both synthetic and real datasets.
Synthetic data: The numeric comparisons on synthetic data can be seen in Table 1. Our proposed method surpasses the recent SOTA method RDLUF-MixS2 according to average PSNR (+0.59 dB) and SSIM (+0.06). Fig. 5 shows the visual reconstruction results. Three wavelengths including striking colors in RGB reference red, yellow, and green are selected to compare. The golden box part in the reference was chosen to calculate and compare the wavelength accuracy. The accuracy metric is the correlation coefficient with the ground truth of the chosen region, i.e. the ‘Corr’ in the curves. According to the ‘Corr’, our method (0.9995) has a more accurate wavelength curve than others. The zoomed part in the figure also demonstrates that our method has clearer edges on the hat than others. In Table. 2, we list the performance, parameter number, FLOPs, inference time, and training time of our method and other recent unfolding methods. ‘Infer. Time’ denotes the total inference time of each method dealing with 10 synthetic test scenes. ‘Training Time’ is under 300 total epochs setting. The table illustrates that our methods exhibit superior performance, the lowest FLOPs, along with reasonable parameter counts and times. Real data: Two scenes of real SD-CASSI measurement reconstruction results are shown in Fig. 6, and two obvious color regions in RGB references are selected to compare. Our method shows less noise and artefacts on the plastic toy surface.
@@ 6 Ablation Study
In this ablation study, we train our model on the synthetic training data with models with 3 unfolding stages. The results are summarized in Table 3. ‘Inaccurate guidance’ denotes that we only use measurement as the latent encoder input instead of clean HSIs. ‘w/o prior flow’ denotes that using simple MLP layers instead of CPF. ‘w/o diffusion’ denotes removing diffusion in Phase II. {tabwindow}[0,r, $\bm{z}$ scale ($N$) 4 16 64 256 PSNR 37.99 38.40 38.40 38.41 FLOPs 33.31 33.80 34.49 36.99 ,$\bm{z}$ scale comparisons.] The ablation illustrates that with LDM prior assistance, we can achieve better reconstruction results, and our design of the Trident Transformer successfully aggregates three types of information and effectively compensates for some reconstruction defects. Fig. 2 also visualizes the feature map changes before and after prior enhancement. The enhanced features demonstrate increased concentration on edges and reduced noise. Moreover, without accurate guidance, the LDM will even harm the reconstruction. The method ‘Basic MHSA’ illustrates that ACS-MHSA in CSF has better performance and higher computational efficiency than basic MHSA. The diffusion steps shown in Fig. 2(b) illustrate that 16 steps are enough for good reconstruction results. Table 2 illustrates that inference time is still in a reasonable range even using diffusion 16 steps. For the scale of $\bm{z}$, we compared different values of $N$ in Table 6, and we find that $N=16$ can keep the balance of performance and efficiency. More ablation studies can be seen in SM.
@@ 7 Conclusion
In this paper, we introduce a novel deep unfolding network that leverages prior knowledge from the latent diffusion model for spectral reconstruction. It achieves SOTA performance on both simulated data and real data.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (grant number 62271414), Zhejiang Provincial Outstanding Youth Science Foundation (grant number LR23F010001), Zhejiang “Pioneer” and “Leading Goose” R&D Program (grant number 2024SDXHDX0006, 2024C03182), the Key Project of Westlake Institute for Optoelectronics (grant number 2023GD007), and the 2023 International Sci-tech Cooperation Projects under the purview of the “Innovation Yongjiang 2035” Key R&D Program (grant number 2024Z126).